Controls Meeting 8-Aug-2013
- Minutes from previous meeting 13-Jun-2013
- Update on JInventory database - Serguei P
- Solenoid future plans - Dave, Mark, Scot, Elliott
- PLC softare, archiving, etc.
- computers and networking
- FCAL dark room update - John, Manuel, Dan
- EPICS - Hovanes, Nerses
- Counting House status - Hovanes
- Other controls applications - Dave
- Storing/retrieving PLC parameters
- FDC cooling, target, BCAL cooling, etc.
- Mantis task tracker
- Midas elog
- Future topics
- Experiment Control Supervisor - Vardan, Elliott
9:30AM Thurs 8-Aug-2013 CC A110
9:30am Thurs 22-Aug-2013
Present: Elliott W, Serguei P, Dave B, Scot S, Mark S, Vanik K.
Serguei presented an update on progress on the JInventory system, see his slides for complete details.
Elliott gave a brief summary on the completion of the solenoid test. The refrigerator finally behaved for long enough to test and map the magnet at up to 1350A (below the target of 1500A). The refrigerator will be repaired, it's not clear what current we will be able to run at with this refrigerator. Many repairs, upgrades and improvements are planned for the solenoid and its control system over the next year (George is maintaining a complete list).
Vanik reported that the goniometer gui is done. The goniometer will soon be tested in the FEL for accuracy (rotation needs 5 microRadian precision!). Also the CDC HV gui is done (by Nerses, I think).
Elliott reported that installation of computers and network cabling in the control room is almost complete. These are needed for the upcoming online data challenge.
Dave reported he'll be working on the FCAL PLC soon, once he's out of Blue Crab. The CDC and FDC will also use CompactLogix systems.